OpenStudy (anonymous):

Problem: Jeanie has 3/4 yard piece of ribbon. She needs one 3/8 yard piece and one 1/2 yard piece. Can she cut the piece of ribbon into the two smaller pieces? Explain!

OpenStudy (anonymous):

No way! Convert 3/4 and 1/2 to have a denominator of 8, and then just add the numerators of the ribbon sizes she want to have.

OpenStudy (anonymous):

Try it...

OpenStudy (anonymous):

She has 3/4 She needs 3/8 + 1/2 If what she needs it greater than what she has, she can't do it.

OpenStudy (anonymous):

She is 1/8 yard short of fabric. Here is what she has minus what she needs. since it is negative, she doesn't have enough fabric.\[\frac{3}{4}*\frac{2}{2}-\left( \frac{3}{8}+\frac{1}{2}*\frac{4}{4} \right)=\frac{6}{8}-\frac{3+4}{8}=-\frac{1}{8}\]

OpenStudy (anonymous):

thanks! your great!